# A Multi-Modal Explainability Approach for Human-Aware Robots in Multi-Party Conversation

### v1

@@ Abstract
The addressee estimation (understanding to whom somebody is talking) is a fundamental task for human activity recognition in multi-party conversation scenarios. Specifically, in the field of human-robot interaction, it becomes even more crucial to enable social robots to participate in such interactive contexts. However, it is usually implemented as a binary classification task, restricting the robot’s capability to estimate whether it was addressed and limiting its interactive skills. For a social robot to gain the trust of humans, it is also important to manifest a certain level of transparency and explainability. Explainable artificial intelligence thus plays a significant role in the current machine learning applications and models, to provide explanations for their decisions besides excellent performance. In our work, we a) present an addressee estimation model with improved performance in comparison with the previous SOTA; b) further modify this model to include inherently explainable attention-based segments; c) implement the explainable addressee estimation as part of a modular cognitive architecture for multi-party conversation in an iCub robot; d) propose several ways to incorporate explainability and transparency in the aforementioned architecture; and e) perform a pilot user study to analyze the effect of various explanations on how human participants perceive the robot.
@@ 1 Introduction
As the hallmark of human-centered technology, the development of interactive robots is increasingly focused on fostering human trust in artificial systems. Therefore, the accuracy and robustness of the performances are essential but not the only indicators of the system’s reliability. Explainability and transparency are two other critical aspects of designing and evaluating reliable interactive robots (Wortham et al.,, 2016). Both allude to the understandability of the system: transparency refers more to the visibility of underlying processes leading to a reduction of ambiguity regarding a behavior (Selkowitz et al.,, 2017), whereas explainability is related to the capability of a system to exhibit the reasons behind its outputs, decisions or behaviors (Ciatto et al.,, 2020; Miller,, 2019). These two qualities are desirable from a dual perspective. In the eyes of developers, who need deep comprehension of the robot to design and assess its functioning, and from the point of view of users, who should be able to intuitively interact with the artificial system (Sciutti et al.,, 2018).
In the context of Human-Robot Interaction (HRI), communication is a specific type of interaction. Following Jakobson, (1981), communication involves the exchange of messages between an addresser (who sends the message) and an addressee (who is entailed to receive it). In the case of spoken messages, the communication is verbal but usually comprises non-verbal elements such as gaze, gestures, poses, etc., which are grasped via vision and are often necessary to contextualize the message properly and to address it to the correct agent (Skantze,, 2021). Even though HRI studies rarely go beyond dyadic interactions, the final goal is often bringing robots to social environments, where they are often required to deal with more than one person and, to achieve this aim, be aware of basic social cues ruling multi-party conversations.
@@
Taking into account explainability in the context of human-activity-aware robots requires considering the concept from different perspectives: not only the generation of explanations within the architecture and models controlling the robot’s behavior but also their communication to and reception by users. Hence, this work seeks to bind together such diverse points of view that cannot be examined separately. Starting from this broader approach and with the final aim to endow a social robot with explainable addressee estimation skills for multi-party conversation, the contribution of this work is divided into two intertwined steps, whose methodology is described in Section 3. Specifically, in Subsection 3.1, we design and train an attention-based neural network to optimize a former AE model (Mazzola et al.,, 2023) while extracting explanations at different stages of the inference. In Subsection 3.2, we deploy the newly developed explainable model in a modular robotic architecture to enable the iCub robot to engage in multi-party conversation, implement a multi-modal system to provide real-time explanations of its behavior, and explore the users’ reception of different modalities of explanation (verbal, embodied, visual) in a user study.
@@ 2.1 Attention-based and explainable neural networks
In the context of machine learning, the development of explainable models is steadily rising (Barredo Arrieta et al.,, 2020). The earliest approaches primarily focused on the inherent form of explainability, i.e., employing a model that is straightforward enough for people to understand its behavior and decisions. This slowly changed when the deep learning models got increasingly stronger, eventually substantially surpassing the accuracy of simple models (Krizhevsky et al.,, 2012). Deep learning models are powerful enough to achieve superhuman performance in certain tasks (He et al.,, 2015), yet they often do not provide any reasoning behind their decisions. Thus, the explainability of deep models started to play a crucial role, mainly in their deployment in critical applications.
Two vastly researched and popular methods of interpretability of deep learning are SHAP (Lundberg and Lee,, 2017) and LIME (Ribeiro et al.,, 2016). These and many similar methods work by training a surrogate, often much simpler model, approximating the black-box model to produce similar output while being inherently explainable. On the other hand, there is often a need to generate precise explanations for the image classification domain. Therefore, saliency maps, i.e., the importance of image regions for predictions, are often investigated (Simonyan et al.,, 2014; Zhou et al.,, 2016; Selvaraju et al.,, 2017). Using these methods, it is possible to peek inside the black-box model processing an image. A downside of similar approaches is their dependence on a specific type of architecture (smooth gradients, convolutions, etc.); otherwise, they are often unable to produce satisfying results.
A branch of research, currently setting the SOTA in the majority of tasks, was initiated by designing the transformer architecture (Vaswani et al.,, 2017) followed by its adaptation for the image domain (Dosovitskiy et al.,, 2021). Thanks to their attention mechanism, transformers are also often considered more interpretable (Kashefi et al.,, 2023), as extracting the attention weights during the forward pass is possible. Thus, in this work, we leverage the idea of attention in general (Shen et al.,, 2017; Niu et al.,, 2021) and the many ways of implementing it in the specific downstream task of AE where, to our knowledge, explainability was never taken into account before.
@@ 2.2 Multi-party conversation in HRI
The management of multi-party conversations requires robots to be endowed with human activity recognition capabilities; this is even more challenging when they have to deal with multiple humans at the same time. Several tasks need to be solved to this aim, not only sound detection and natural language understanding, which are essential to receiving the message. Speaker recognition and diarization, turn-taking, and addressee estimation are crucial problems that need to be tackled to assess, beyond the “what" of the message, the “who" and the “to whom" of each utterance (Gu et al.,, 2022). Endowed with multiple sensors, robots can solve problems related to multi-party conversations with a multi-modal approach to recognize the scene and human intentions via audio, vision, and, additionally, interpreting information coming from the conversation context (Bilac et al.,, 2017; Dhaussy et al.,, 2023; Bae and Bennett,, 2023; Addlesee et al.,, 2024).
@@
Differently from dyadic scenarios, multi-party conversations present an additional problem once the speaker yields the turn: who is entitled to take it? If the conversation must proceed, a correct estimation of the utterance’s addressee(s) usually implies who should take the turn. In the majority of cases, works tackling AE during the interaction with artificial agents designed rule-based algorithms (Richter et al.,, 2016), machine learning models (van Turnhout et al.,, 2005; Huang et al.,, 2011; Sheikhi et al.,, 2013), deep neural networks (Mazzola et al.,, 2023; Tesema et al.,, 2023) or large language model (LLM) based techniques (Addlesee et al.,, 2024) grounded on multiple modalities and features in order to cope with the ambiguity and unpredictability of human behaviors in real-time interactions. Keywords uttered by speakers, their gaze, pose, para-verbal cues and contextual information have all been demonstrated useful to the purpose of AE (for a review, see Skantze, (2021)). AE may improve robots’ conversational abilities by providing information not only about when intervening but how to do so. However, most current models predict only whether the robot was addressed in a binary way (van Turnhout et al.,, 2005; Huang et al.,, 2011; Sheikhi et al.,, 2013; Tesema et al.,, 2023; Addlesee et al.,, 2024). But binary estimation does not identify the addressee (if it is other than the robot) and, therefore, is insufficient for the robot to effectively engage in conversations with more than two humans or without pre-determined knowledge about other users’ presence.
@@
To resolve this limitation, Mazzola et al., (2023) adopted a deep-learning approach to estimate the direction of the addressee from the robot’s perspective, training the model on HRI data collected with a Nao robot (Jayagopi et al.,, 2012). The same model was then ported and tested with a pilot experiment on the iCub platform (Mazzola et al.,, 2024), but not yet implemented in an architecture for multi-party conversation. Such implementation is one of the goals of the present study: after optimizing the approach of Mazzola et al., (2023), we incorporate the new explainable AE model into a modular architecture with additional components (e.g., spatial memory and action manager) to make the robot find and identify the addressee over its limited field of view.
@@ 2.3 Explainability in HRI
Explainable artificial intelligence (XAI) has predominantly been explored in the human-computer interaction field (Lai et al.,, 2021; Gambino and Liu,, 2022) and even though the number of works concerning robots’ explainability is increasing, there are still a few studies about XAI within the HRI context (De Graaf and Malle,, 2017). Robots introduce an additional layer of complexity to the explainability problem compared to virtual agents due to their embodiment and the broader spectrum of interaction modalities they offer (Setchi et al.,, 2020).
Automatic explanation generation with robots has been investigated in several interaction contexts, such as planning (Chakraborti et al.,, 2017) and human-robot collaboration (Matarese et al., 2023a, ). For instance, Chakraborti et al., (2017) generated explanations while trying to resolve the discrepancies between the robot and human’s internal models. Diversely, Tabrez et al., (2019) tackled the problem by focusing on users’ task understanding to detect incomplete or incorrect beliefs about the robot’s functioning. Matarese et al., 2023a focused on explainable robots’ influence when providing explanations that consider the human-robot common ground, also regarding people’s personality traits (Matarese et al., 2023b, ).
Visual explanations with robots have been proposed for several purposes, such as navigation (Maruyama et al.,, 2022; Halilovic and Lindner,, 2023). However, more importantly for the scope of this work, Zhu et al., (2022) proposed a multi-modal explanation framework that coupled visual-based with verbal-based explanations to explain facial emotion recognition. Moreover, Sobrín-Hidalgo et al., (2024) presented a preliminary study proposing a vision-language model that allows the robot to generate explanations combining data from its logs and the images it captures.
In recent years, the HRI community has shown a growing interest in verbal explanations that is destined to grow further, given the spreading of LLMs. Stange et al., (2022) designed and developed a dialogical model for explanations in HRI. With their model, the robot can reply to human users’ requests with explanations referring to its internal state. The authors stressed the iterative nature of their model in managing the explanatory processes as dialogues. Task understanding has also been investigated from a dialogical perspective with a focus on the role of negation in human-robot explanatory exchanges (Groß et al.,, 2023). Moreover, to allow artificial agents to adapt their explanations to their partners’ understanding, Robrecht and Kopp, (2023) implemented a linguistic explainer model that constructs and employs a partner model.
In the context of the multi-party conversation, explainability has been investigated for sentiment analysis of social media dialogues (Sinha et al.,, 2021) and emotion recognition with multi-modal attentive learning (Arumugam et al.,, 2022). However, to the best of our knowledge, there is no approach to provide real-time explainable and transparent solutions for robot’s behavior in multi-party interaction.
@@ 3.1 Design of the attention-based explainable AE model
The development of our explainable AE model consists of two steps: in the first step, we focus on enhancing the classification accuracy with respect to the previous SOTA in the same task (Mazzola et al.,, 2023), which represents our baseline (see Paragraph 3.1.1). This way, we obtain a first model, which we refer to as Improved Addressee Estimation (IAE) model. In the second step, we modify this model with inherently explainable modules based on attention (see Paragraph 3.1.2) to extract additional information during addressee estimation. We refer to the second model as Explainable Addressee Estimation (XAE) model.
@@ 3.1.1 Improved Addressee Estimation model
To consistently compare with the baseline (Mazzola et al.,, 2023), we only use the conversation parts in which ROBOT, RIGHT, or LEFT is the target. We also follow the same data pre-processing. The view from the robot’s camera is split into two parallel data streams (face images and body-pose vectors), which serve as an input to the network and later are merged to form a combined representation.
We perform a hyper-parameter search (see details in the Appendix) to explore the achievable prediction accuracy on the given task. We always keep one of the 10 interactions included in the dataset for testing and perform 9-fold cross-validation on the remaining 9 interactions. This cross-validation performance (weighted according to the number of sequences in individual interactions) is being optimized.
Our chosen IAE model improves the current SOTA while significantly reducing the number of trainable parameters $\approx$135 folds (from 91,706,749 to 677,623). This is achieved mainly by reducing the number of output neurons from the convolutional neural network (CNN) processing the facial information. Furthermore, we replace the convolution on body-pose vectors with fully connected layers. The output dimensionality is chosen to make the length of the outputs from face and pose models similar, allowing for more sophisticated data-fusion methods.
3.1.2 XAE model: incorporating attention

### v2

@@ Abstract
The addressee estimation (understanding to whom somebody is talking) is a fundamental task for human activity recognition in multi-party conversation scenarios. Specifically, in the field of human-robot interaction, it becomes even more crucial to enable social robots to participate in such interactive contexts. However, it is usually implemented as a binary classification task, restricting the robot’s capability to estimate whether it was addressed or not, which limits its interactive skills. For a social robot to gain the trust of humans, it is also important to manifest a certain level of transparency and explainability. Explainable artificial intelligence thus plays a significant role in the current machine learning applications and models, to provide explanations for their decisions besides excellent performance. In our work, we a) present an addressee estimation model with improved performance in comparison with the previous state-of-the-art; b) further modify this model to include inherently explainable attention-based segments; c) implement the explainable addressee estimation as part of a modular cognitive architecture for multi-party conversation in an iCub robot; d) validate the real-time performance of the explainable model in multi-party human-robot interaction; e) propose several ways to incorporate explainability and transparency in the aforementioned architecture; and f) perform an online user study to analyze the effect of various explanations on how human participants perceive the robot.
@@
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
As the hallmark of human-centered technology, the development of interactive robots is increasingly focused on fostering human trust in artificial systems. Therefore, the accuracy and robustness of the performances are essential but not the only indicators of the system’s reliability. Explainability and transparency are two other critical aspects of designing and evaluating reliable interactive robots (Wortham et al.,, 2016). Both allude to the understandability of the system: transparency predominantly refers to the visibility of underlying processes leading to a reduction of ambiguity regarding a behavior (Selkowitz et al.,, 2017), whereas explainability is related to the capability of a system to exhibit the reasons behind its outputs, decisions or behaviors (Ciatto et al.,, 2020; Miller,, 2019). These two qualities are desirable from a dual perspective. In the eyes of developers, who need deep comprehension of the robot to design and assess its functioning, and from the point of view of users, who should be able to intuitively interact with the artificial system (Sciutti et al.,, 2018).
In the context of Human-Robot Interaction (HRI), communication is a specific type of interaction. Following Jakobson, (1981), communication involves the exchange of messages between an addresser (who sends the message) and an addressee (who is entailed to receive it). In the case of spoken messages, the communication is verbal but usually comprises non-verbal elements such as gaze, gestures, poses, etc., which are grasped via vision and are often necessary to properly contextualize the message and to address it to the correct agent (Skantze,, 2021). Even though HRI studies rarely go beyond dyadic interactions, the final goal is often bringing robots to social environments, where they are often required to deal with more than one person and, to achieve this aim, be aware of basic social cues ruling multi-party conversations.
@@
Taking into account explainability in the context of human-activity-aware robots requires considering the concept from different perspectives: not only the generation of explanations within the architecture and models controlling the robot’s behavior but also their communication to and reception by users. Hence, this work seeks to bind together such diverse points of view that cannot be examined separately. Starting from this broader approach and with the final aim to endow a social robot with explainable addressee estimation skills for multi-party conversation, we address several key gaps in the field of addressee estimation (AE) for human-robot interaction (HRI). First, we introduce the concept of explainability into AE, leveraging attention-based mechanisms, which, to our knowledge, have not been previously explored in this context. Second, we tackle the limitations of existing AE approaches, which typically only determine if the robot is being addressed, by enhancing the system to identify specific addressees in multi-party conversations. Building on the work of Mazzola et al., (2023), we optimize their addressee direction estimation model and integrate it into a modular architecture for multi-party interaction, incorporating additional components to enable the robot not only to understand basic communication dynamics but also to leverage communication cues and infer the presence of new people in the environment. Finally, while explainability has been studied in related areas such as sentiment analysis and emotion recognition, we provide the first real-time explainable solution for robotic behavior in multi-party conversations, offering both enhanced transparency and practical utility for HRI scenarios.
The contribution of this work is divided into two intertwined steps, whose methodology is described in Section 3. Specifically, in Subsection 3.1, we design and train an attention-based neural network to optimize a former AE model (Mazzola et al.,, 2023) while extracting explanations at different stages of the inference. In Subsection 3.2, we deploy the newly developed explainable model in a modular robotic architecture to enable the iCub robot to engage in multi-party conversation, implement a multi-modal system to provide real-time explanations of its behavior, and explore the users’ reception of different modalities of explanation (verbal, embodied, graphical) in a user study.
@@ 2.1 Attention-based and explainable neural networks
Explainable models are becoming increasingly important in machine learning (Barredo Arrieta et al.,, 2020). Early efforts focused on using simple models to ensure inherent interpretability. However, as deep learning models improved and outperformed simpler ones (Krizhevsky et al.,, 2012), achieving even superhuman performance in some tasks (He et al.,, 2015), their lack of transparency became a significant concern. This has made explainability essential, particularly for deploying deep models in critical applications.
Two widely researched and popular methods for deep learning interpretability are SHAP (Lundberg and Lee,, 2017) and LIME (Ribeiro et al.,, 2016). These, along with similar approaches, train a simpler surrogate model to approximate the black-box model’s outputs, making it inherently more explainable.
Latent representations of inputs are often analyzed to understand model behavior. However, their high dimensionality often makes visualization and explanation challenging. To achieve this, dimensionality-reduction techniques can project the complex data into 1D, 2D, or 3D space, enabling clearer data visualization. Principal Component Analysis (PCA), t-distributed Stochastic Neighbor Embedding (t-SNE), and Uniform Manifold Approximation and Projection (UMAP) are commonly used to reduce the space dimensionality, each having its advantages and disadvantages (van der Maaten and Hinton,, 2008; McInnes et al.,, 2018).
On the other hand, there is often a need to generate precise explanations for the image classification domain. For this, saliency maps, i.e., the importance of image regions for predictions, are often investigated (Simonyan et al.,, 2014; Zhou et al.,, 2016; Selvaraju et al.,, 2017). Using these methods, it is possible to peek inside the black-box model processing an image. A downside of these approaches is their dependence on a specific type of architecture (smooth gradients, convolutions, etc.); otherwise, they are often unable to produce satisfying results.
A branch of research, currently setting the state-of-the-art (SOTA) in the majority of tasks, was initiated by designing the transformer architecture (Vaswani et al.,, 2017) followed by its adaptation for the image domain (Dosovitskiy et al.,, 2021). Thanks to their attention mechanism, transformers are also often considered more interpretable (Kashefi et al.,, 2023), as one can extract the attention weights during the forward pass.
Thus, in this work, we leverage the idea of attention in general (Shen et al.,, 2017; Niu et al.,, 2021) and the many ways of implementing it in the specific downstream task of AE where, to our knowledge, explainability was never taken into account before.
@@ 2.2 Multi-party conversation in HRI
The management of multi-party conversations requires robots to be endowed with human activity recognition capabilities. Several tasks need to be solved to this aim, not only sound detection and natural language understanding, which are essential to receiving the message. Speaker recognition and diarization, turn-taking, and addressee estimation are crucial problems that need to be tackled to assess beyond the “what" of the message, the “who" and the “to whom" of each utterance (Gu et al.,, 2022). Endowed with multiple sensors, robots can solve multi-party conversation problems with a multi-modal approach to recognize the scene and human intentions via audio, vision, and interpreting information coming from the conversation context (Bilac et al.,, 2017; Dhaussy et al.,, 2023; Bae and Bennett,, 2023; Addlesee et al.,, 2024).
Unlike dyadic scenarios, multi-party conversations introduce the challenge of determining who should take the turn when the speaker yields. For the conversation to continue, correctly identifying the addressee(s) typically indicates the next speaker. Keywords, gaze, pose, para-verbal cues, and contextual information have all proven useful for Addressee Estimation (AE) (Skantze,, 2021). AE enhances robots’ conversational abilities by informing both when to intervene and how to do so.
@@
In the majority of cases, works tackling AE during the interaction with artificial agents designed rule-based algorithms (Richter et al.,, 2016), machine learning models (van Turnhout et al.,, 2005; Huang et al.,, 2011; Sheikhi et al.,, 2013), deep neural networks (Mazzola et al.,, 2023; Tesema et al.,, 2023) or large language model (LLM) based techniques (Addlesee et al.,, 2024), grounded on multiple modalities and features to cope with the ambiguity and unpredictability of human behaviors in real-time interactions. However, most current models predict only whether the robot was addressed in a binary way (van Turnhout et al.,, 2005; Huang et al.,, 2011; Sheikhi et al.,, 2013; Tesema et al.,, 2023; Addlesee et al.,, 2024). This approach is unsuitable for multi-party interactions; therefore, it is insufficient for the robot to effectively engage in conversations with more than two humans without pre-determined knowledge about other users’ presence.
@@
To resolve this limitation, Mazzola et al., (2023) adopted a deep-learning approach to estimate the direction of the addressee from the robot’s perspective, training the model on HRI data collected with a Nao robot (Jayagopi et al.,, 2012). The same model was then ported and tested with a pilot experiment on the iCub platform (Mazzola et al.,, 2024), but not yet implemented in an architecture for multi-party conversation. After optimizing the approach of Mazzola et al., (2023), we implemented the new explainable AE model into a more complex modular architecture to allow the robot to identify the addressee outside its field of view.
@@ 2.3 Explainability in HRI
Explainable Artificial Intelligence (XAI) has predominantly been explored in the human-computer interaction field (Lai et al.,, 2021; Gambino and Liu,, 2022) but there are still a few studies about XAI within the HRI context (De Graaf and Malle,, 2017). Robots introduce additional layers of complexity to the explainability problem than virtual agents due to their embodiment and the higher number of interaction modalities they can handle (Setchi et al.,, 2020).
Regarding robot planning, Chakraborti et al., (2017) generated explanations while trying to resolve the discrepancies between the robot and human’s internal models. Diversely, Tabrez et al., (2019) tackled the problem by focusing on users’ task understanding to detect incomplete or incorrect beliefs about the robot’s functioning. Instead, concerning human-robot collaboration, Matarese et al., 2023a focused on explainable robots’ influence when providing explanations that consider their common ground, also regarding people’s personality traits (Matarese et al., 2023b, ).
Visual explanations have also been explored in robotics (Maruyama et al.,, 2022; Halilovic and Lindner,, 2023). Notably, Zhu et al., (2022) introduced a multi-modal framework combining visual and verbal explanations for facial emotion recognition. Additionally, Sobrín-Hidalgo et al., (2024) presented a preliminary study on a vision-language model enabling robots to generate explanations by integrating log and video data.
In recent years, the HRI community has shown a growing interest in verbal explanations, destined to grow further because of the spreading of LLMs. For example, Stange et al., (2022) designed and developed a dialogical model for explanations in HRI, which allows robots to reply to human users’ requests and explain their internal state. The authors stressed the iterative nature of their model in managing the explanatory processes as dialogues. Task understanding has also been investigated from a dialogical perspective focusing on the role of negation in human-robot explanatory exchanges (Groß et al.,, 2023). Moreover, to allow artificial agents to adapt their explanations to their partners’ understanding, Robrecht and Kopp, (2023) implemented a linguistic explainer model that constructs and employs a partner model.
In the context of multi-party conversation, explainability has been investigated for sentiment analysis of social media dialogues (Sinha et al.,, 2021) and emotion recognition with multi-modal attentive learning (Arumugam et al.,, 2022). However, to our knowledge, there is no approach to provide real-time explainable and transparent solutions for robot behavior in multi-party interaction.
@@ 3.1 Design of the attention-based explainable AE model
The development of our explainable AE model consists of two steps: in the first step, we focus on enhancing the classification accuracy with respect to the previous SOTA in the same task (Mazzola et al.,, 2023), which represents our baseline (see Paragraph 3.1.1). This way, we obtain a first model, which we refer to as the Improved Addressee Estimation (IAE) model. In the second step, we modify this model using inherently explainable modules based on attention (see Paragraph 3.1.2) to extract additional information during addressee estimation. We refer to the second model as Explainable Addressee Estimation (XAE) model [^1] The implementation can be found here: https://github.com/iveta331/Explainable-Addressee-Estimation.
@@ 3.1.1 Improved Addressee Estimation model
To consistently compare with the baseline (Mazzola et al.,, 2023), we only use the conversation parts in which ROBOT, RIGHT, or LEFT is the target. We also follow the same data pre-processing and data augmentation. To fully leverage the Vernissage dataset, we further augment the input images using commonly recognized techniques. It involves diversifying the lightning conditions (slight changes in brightness, contrast, saturation, and hue), applying spatial changes (rotation and crop), and blurring the images. The exact ranges for each of the data augmentation methods are detailed in Tab. 1. The view from the robot’s camera is split into two parallel data streams (face images and body-pose vectors), which serve as input to the network and later are merged to form a combined representation.
The model from (Mazzola et al.,, 2023) is further modified by reducing the number of output neurons from the convolutional neural network (CNN) processing the facial information. Also, we replace the convolution on body-pose vectors with fully connected layers. The output dimensionality is chosen to make the length of the outputs from face and pose models similar, allowing for more sophisticated data-fusion methods. Thanks to these modifications, our IAE model improves the current SOTA while significantly reducing the number of trainable parameters $\approx$135 folds (from 91,706,749 to 677,623).
3.1.2 Explainable addressee estimation model